OpenStudy (anonymous):

solve -5/9=-2v+1/3 for v. 5/6 1/9 4/9 -1/18

OpenStudy (anonymous):

v = 4/9

OpenStudy (anonymous):

-5/9=-2v+1/3 -2v = -5/9 - 1/3 -2v = -5/9 - 3/9 -2v = - 8/9 v = 8/9 * 1/2 v = 4/9